The sum of a number and twice its multiplicative inverse (reciprocal) is 3. Find the number.
Answer by mananth(16949)

x be the number
reciprocal = 1/x
2 times reciprocal = 2/x
...
x+2/x = 3
...
(x^2+2)/x=3
(x^2+2)=3x
x^2-3x+2=0
x^2-2x-x+2=0
x(x-2)-1(x-2)=0
(x-2)(x-1)=0
x = 1 OR 2
...
m.ananth@hotmail.ca